OpenStudy (cutiecomittee123):

Solve for x 500=25e^2x

OpenStudy (mle):

divide both sides by 25

OpenStudy (mle):

so you get 20 = e^(2x) right? so ln base e of 20 = 2x then divide both sides by 2 to get x alone

OpenStudy (cutiecomittee123):

okay

OpenStudy (cutiecomittee123):

x=10

OpenStudy (anonymous):

Divide both sides by 25x. 25e2/x25x=500/25x a2=20/x a=20/x or a=−20/x Answer: a=20/x or a=−20/x

OpenStudy (cutiecomittee123):

wouldnt it be 1n(20)/2

OpenStudy (cutiecomittee123):

because 2x is an exponent

OpenStudy (anonymous):

Sorry, I read it as "a" squared time x let me fix it.

OpenStudy (mle):

yes cutie, I believe that is correct ^_^